A study followed a cohort of students entering a
business/economics program.17 All students followed a common track during the first three semesters and then
chose to specialize in either business or economics. Through a series of surveys, the researchers were able to
classify roughly 50% of the students as either peer driven (ignored abilities and chose major to follow peers) or
ability driven (ignored peers and chose major based on ability). When looking at entry wages after graduation, the
researchers conclude that a peer-driven student can expect an average wage that is 13% less than that of an
ability-driven student. The report states that the significance level is P = 0.09. Can you be confident of the
researchers’ conclusion statement regarding the wage decrease? Explain your answer.

PLZ ANYONE CAN HELP ME OUT WITH THIS...
Answer #1

The null hypothesis of this test (H0) is P = S1 / S2 = (1 - 0.13) i.e. 0.87 vs. the alternative hypothesis (HA): p > 0.87

Where S1 and S2 are the average entry wages for the peer-driven and ability-drive students respectively.

The result of the test shows a p-value = 0.09 which is more than alpha = 0.05

This means that the null hypothesis cannot be rejected at a 5% significance level. So, at this significance level, the peer-driven student can expect an average wage less than that of an ability-driven student by 13% can be supported because that was the null hypothesis.